An aqueous solution contains 1 mole of urea and 1kg of water. Kf for water is 1.8 kg/mol. Calculate the amount of ice seperated from the solution at -3.6°C.

Grade:12

1 Answers

Vikas TU
5 years ago
From T = i*Kf*m
i = 1
kf = 1.8 kg/mol
 
T = 0- (-3.6) = 3.6 
3.6 = 1.8*m
m = 2
m’ = moles of urea/mass of solvent in kg => 1/1 = 1
m – m’ = 2-1 = 1
1 = w/18
w = 18 gms of ice has been separated.